# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

### v2

@@ 2.2. Content-based Filtering
Content-based filtering (CBF) (Pazzani and Billsus, 2007) recommends items by analyzing their attributes and comparing them to a user’s preferences or past interactions. This technique creates user profiles based on demographic information or previous interactions and aligns these with detailed item descriptions to personalize recommendations. It addresses the cold start problem effectively by suggesting items based on their features, even without prior user interaction history. While adept at recommending items similar to those a user has previously shown interest in, CBF can sometimes lack diversity in its suggestions. To overcome this, recent advancements leverage machine learning and natural language processing to refine user-profiles and item descriptions, aiming to improve the relevancy and accuracy of recommendations. Despite its challenges, CBF remains a key method for offering personalized recommendations, with continuous enhancements driven by technological progress.
@@ 2.3. Hybrid Recommender Systems
Hybrid recommender systems combine elements from both collaborative filtering and content-based filtering to offer more accurate and diverse recommendations. By integrating these approaches, hybrid systems can leverage the strengths and mitigate the weaknesses of each method. They use a variety of techniques, including combining separate collaborative and content-based models, incorporating features of one method into the other, or building a unified model that simultaneously learns user and item representations from both user interactions and item attributes. This approach allows hybrid systems to provide personalized recommendations like content-based methods while also benefiting from the broad recommendation capabilities of collaborative filtering. Hybrid systems are particularly effective in addressing the cold start problem for new users and items, as they can recommend items based on content when user interaction data is sparse. However, the hybrid recommender system is complex and expensive to implement (Su and Khoshgoftaar, 2009).
@@ 3. Why adding review into recommender systems
The ultimate goal of a recommender system is to suggest items that match user’s interests. Traditional recommender systems learn the user’s preferences based on historical rating data. However, these rating-based systems encounter significant challenges, including data sparsity, a lack of explainability, and reduced accuracy. Incorporating additional forms of information, such as text and images, is advantageous when available. Reviews, in particular, play a crucial role in understanding user preferences and item characteristics. User reviews, being free-form text, allow users to express their opinions freely in any direction. Utilizing reviews helps address these critical issues in recommender systems effectively.
@@
Addressing Data Sparsity: Data sparsity poses a significant challenge in recommender systems, particularly in real-world scenarios where rating data is scarce, often accompanied by few reviews. Learning user preferences and item features from such sparse data is challenging. Reviews play a crucial role in addressing this challenge. Through textual feedback, users offer detailed opinions, making it valuable for extracting features to understand user preferences. Various existing recommender systems demonstrate the effectiveness of leveraging reviews in mitigating data sparsity issues (Zheng et al., 2017; Tan et al., 2016).
@@
Improving Accuracy: Improving accuracy is another challenging task for any recommender system. Recommendation accuracy largely depends on learning user and item representation. Since reviews contain fine-grained opinions, utilizing them for feature learning can be beneficial. There has been a significant amount of work available that utilizes review to improve accuracy (Kim et al., 2016; Zhang et al., 2017; Tan et al., 2008; Seo et al., 2017). Different NLP techniques are used to extract the features from user reviews and item reviews, followed by an interaction technique to predict the ratings. Some paper utilizes both review and rating for better learning the user and item features as well as improving recommender accuracy (McAuley and Leskovec, 2013).
@@
Enhancing Explainability: Explainability is a significant issue in recommender systems. A recommender system needs to be able to provide a rationale for its recommendations to a user. Traditional rating-based recommender system lacks interpretability because they solely rely on latent factor learning. Latent factors represent the hidden aspects that influence user preferences, but they usually cannot explain why a user gave a specific rating to an item. Reviews can be analyzed to identify the aspects or features that motivate the users to provide specific ratings. Many existing models have demonstrated that utilizing reviews can improve explainability (McAuley and Leskovec, 2013; Seo et al., 2017).
@@ 4. Review-based recommender systems
The three main components of recommender systems are user/item representation learning, user-item feature interaction mechanism, and overall rating or preference prediction. Traditional recommender systems take user $ID$ and item $ID$ as input and learn their features using latent factor models (Koren et al., 2009). However, these static feature learning methods suffer from data sparsity and scalability problems. It is often beneficial to use as much as information possible including the user review text(Chin et al., 2018), images (Yilma and Leiva, 2023), numerical rating score(Cheng et al., 2018a), video (Mu and Wu, 2023), and audio (Bogdanov et al., 2013). Many e-commerce websites such as Amazon.com  [^1] https://www.amazon.com/, Yelp.com [^2] https://www.yelp.com/, and TripAdvisor.com [^3] https://www.tripadvisor.com/ collect user feedback as textual reviews in addition to numerical ratings. User reviews are considered as a rich source of feedback for learning user preferences or item attributes. Because users are free to express their opinions, it is assumed that the users’ fine-grained preferences are available in the review. These reviews are utilized in various ways to solve problems in traditional rating-based recommender systems, e.g. data sparsity problems, and explainability problems.
Figure 2. Sample Hotel Review Highlighting Key Aspects
Figure 3. The detailed workflow of the review-based recommender systems. The top panel represents the Review-Only Recommender Systems, and the bottom panel represents the Rating and Review-Integrated Recommender Systems
We categorize the review-based recommender systems into two categories: 1) General review-based method, and 2) Aspect-based method. General review-based methods (Musat et al., 2013; Liu et al., 2019a; Zheng et al., 2017) learn overall user preferences and item characteristics from the review without considering user sentiment on any particular topic or aspects of the items. In contrast, aspect-based methods (Chin et al., 2018) rely on estimating fine-grained sentiment on individual aspects of items. The details for each of the categories are described in the subsequent subsections.
4.1. General Review-based Recommender Systems
To illustrate this category, suppose a user set $U=\{u_{1},u_{2},...,u_{|U|}\}$, item set $I=\{i_{1},i_{2},...,i_{|I|}\}$, review set $E=\{e_{1},e_{2},...,e_{|E|}\}$, and an overall rating set $R=\{r_{1},r_{2},...,r_{|R|}\}$.
In a traditional recommender system, user $u^{\prime}s$ preference $\hat{r}$ towards item $i$ is estimated as
@@ 4.1. General Review-based Recommender Systems
where $P_{u}$ and $Q_{i}$ are the latent feature vector of user $u$ and item $i$ learned using latent factor models, e.g. matrix factorization, and $b$ is the bias term.
The review-based system focuses on integrating review features with rating features to estimate the overall rating. Two different techniques are used to learn user and item representation from reviews: 1) learning user and item representation from review only, and 2) utilizing both reviews and ratings. The former focuses on creating user and item documents to learn user and item representation while the latter emphasizes integrating features from both reviews and ratings. Let’s consider the user and item representation from reviews are $X_{u}$ and $X_{i}$ respectively. Then the equation (1) for the above-mentioned two methods can be rewritten as
@@
and
(3) $\hat{r}=concat(P_{u},X_{u})^{\mathsf{T}}concat(Q_{i},X_{i})+b$
Although the general review-based recommender system demonstrates performance improvement over the rating-only recommender system, it often lacks explainability. This is because the general recommender system focuses on learning the overall preferences for items. To make the system more trustable, it is important to consider the detailed analysis of the review and generate the explanation when recommending particular items.
4.2. Aspect-based Recommender Systems
The aspect-based recommender system aims to understand user’s fine-grained opinions on items. It is done by extracting or learning latent representations of aspects from the reviews. Consider an aspect set $A=\{a_{1},a_{2},...,a_{|A|}\}$ where $|A|$ is the number of aspects and the aspect-based representation of user and item are $X_{U}^{A}$ and $X_{I}^{A}$ respectively. The equation (2) and (3) can be reformulated as
@@ 4.2. Aspect-based Recommender Systems
and
(5) $\hat{r}=concat(P_{u},X_{u}^{A})^{\mathsf{T}}concat(Q_{i},X_{i}^{A})+b$
where $X_{u}^{A}$ and $X_{i}^{A}$ are aspect level representation of users and items respectively. The publications of different categories of review-based papers are listed in Table 1.
The term “aspect” refers to the specific attributes of products. Naturally, a user might not be interested in all aspects of a product but one or more of them. Furthermore, the significance of an aspect to a user can differ from one item to another. Hence, it is vital to capture the aspects and their relevance to users. Aspect-aware recommender models leverage advanced NLP techniques such as topic modeling, deep neural networks, and attention mechanisms to dynamically model user preferences. In Figure 2 “Location”, “service”, and “cost” represent the aspects.
@@
General Review-based (Wei et al., 2024; Chiang et al., 2023; Yang et al., 2023; Xv et al., 2022; Choi et al., 2022; Shuai et al., 2022; Luo et al., 2021; Tan et al., 2021; Xiong et al., 2021; Wang et al., 2021b; Liu et al., 2021d; Gao et al., 2020b; Peña et al., 2020; Dong et al., 2020; Ni et al., 2019; Xia et al., 2019; Wu et al., 2019c; Liu et al., 2019a, b; Rafailidis and Crestani, 2019; Alexandridis et al., 2019; Deng et al., 2018; Wang et al., 2018; Hyun et al., 2018; Chen et al., 2018a; Lu et al., 2018; Zhang et al., 2017; Seo et al., 2017; Catherine and Cohen, 2017; Zheng et al., 2017; He et al., 2015; Almahairi et al., 2015; Liu et al., 2022a; Zhang et al., 2021; Lyu et al., 2021; Wu et al., 2019b; Chen et al., 2019; Mei et al., 2018; Tan et al., 2016; Kim et al., 2016; Wei et al., 2023; Xi et al., 2021; Sayyadpour and Nazarizadeh, 2023; You et al., 2023; Ferdiana et al., 2022; Huang, 2022; Cai et al., 2022; Wu et al., 2021; Du et al., 2021; Liu et al., 2021a, b, c; Shamir et al., 2021; Ray et al., 2020; Huang et al., 2020b; Peng et al., 2020; Khosla et al., 2020; Jin et al., 2020; Suzuki et al., 2019; Cao et al., 2019; Shalom et al., 2019; Li et al., 2019; He et al., 2018; Wu et al., 2019a; Zhang et al., 2016)
Aspect-based (Wang et al., 2023b, a; Yang et al., 2021; Guan et al., 2019; Cheng et al., 2018b, a; Bauman et al., 2017; Li et al., 2015; Shi et al., 2022; Chin et al., 2018; Sun et al., 2021; You et al., 2020; Da’u et al., 2020; Mowlaei and Keyvanpour, 2020; Tewari et al., 2019; Sangeetha et al., 2017; Qiu et al., 2016; Bhojwani et al., 2023; Won et al., 2022; Ozcan et al., 2022; Peng et al., 2022; Cantador et al., 2022; Liu et al., 2022b; V and Shah, 2022; Godakandage et al., 2020; Da’u and Salim, 2019; Wu and Ester, 2015; Safar et al., 2022)
Each of these categories can further be subdivided into two classes: a) Review-Only Recommender Systems, and b) Rating and Review-Integrated Recommender Systems.
Review-Only Recommender Systems: Within this classification, recommender systems utilize reviews exclusively to understand user behavior and item attributes. In this case, user and item reviews are used to construct user and item profiles, resulting in improved representation of user-item interactions. This is illustrated in the upper panel of Figure 3.
Rating and Review-Integrated Recommender Systems: Both the rating and review-based features are bridges to obtain a rich representation of the user and item. The combined strength of rating-based and review-based features is harnessed to improve feature representation. This category is depicted in the lower panel of Figure 3.
5. State-of-the-art Review-aware Recommender Systems
In this section, we briefly present the state-of-the-art review-based recommender systems. Our analysis emphasizes on the methodology, the rationale behind the method, features, datasets, evaluation metrics, and performance. We classify these systems based on the methods used into two main categories: 1) Probabilistic Approach, and 2) Deep Learning-based Approach.
5.1. Probabilistic Approaches

### v3

@@ 2.2. Content-based Filtering
Content-based filtering (CBF) recommends items by analyzing their attributes and comparing them to a user’s preferences or past interactions (Pazzani and Billsus, 2007). This technique creates user profiles based on demographic information or previous interactions, then aligns these with item attributes for personalized recommendations. It addresses the cold start problem effectively by suggesting items based on their features, even without prior user interaction history. While adept at recommending items similar to those a user has previously shown interest in, CBF can sometimes lack diversity in its suggestions. To overcome this, recent advancements leverage machine learning and natural language processing to refine user profiles and item descriptions, aiming to improve the accuracy, relevancy and diversity of recommendations (Yang et al., 2021; Wang et al., 2021b; Lyu et al., 2021). Despite its challenges, CBF remains a key method for personalized recommendations, with continuous enhancements driven by technological progress.
@@ 2.3. Hybrid Recommender Systems
Hybrid recommender systems combine elements from both collaborative filtering and content-based filtering to offer more accurate and diverse recommendations (Lucas et al., 2013). By integrating these approaches, hybrid systems can leverage the strengths and mitigate the weaknesses of each method. They use a variety of techniques, including combining separate collaborative and content-based models (De Campos et al., 2010; Choi et al., 2010), incorporating features of one method into the other (Bedi et al., 2013; Mooney and Roy, 2000), or building a unified model that simultaneously learns user and item representations from both user interactions and item attributes (Çano and Morisio, 2017). These techniques allow hybrid systems to provide personalized recommendations like content-based methods while also benefiting from the broad recommendation capabilities of collaborative filtering (Choi et al., 2010). Hybrid systems are particularly effective in addressing the cold start problem for new users and items, as they can recommend items based on content when user interaction data is sparse (Thorat et al., 2015). For instance, sentiment analysis and multi-criteria review mining can extract valuable insights from user reviews, including specific opinions on product aspects and overall user sentiment. Sentiment-Aware Deep Learning (Hung, 2020) and Aspect-Based Sentiment Analysis (Musto et al., 2017) have shown to enhance recommendation quality, particularly in scenarios with sparse user ratings or new items. However, implementing hybrid recommender systems can be complex and resource-intensive (Su and Khoshgoftaar, 2009).
@@ 3. Why adding review into recommender systems
The ultimate goal of a recommender system is to suggest items that match user’s interests. Traditional recommender systems learn user’s preferences based on historical rating data. However, these rating-based systems encounter significant challenges, including data sparsity, lack of explainability, and reduced accuracy. Incorporating additional forms of information, such as text and images, is advantageous when available. Reviews, in particular, play a crucial role in understanding user preferences and item characteristics. User reviews, being free-form text, allow users to express their opinions in any direction. Utilizing reviews helps addressing these critical issues in recommender systems. Following are some of the key avantages of incorporating reviews.
@@
Addressing Data Sparsity: Data sparsity poses a significant challenge in recommender systems, particularly in real-world scenarios where rating data is scarce. Learning user preferences and item features from such sparse data is challenging. Reviews help address this challenge by providing detailed textual feedback, which is valuable for extracting features to understand user preferences. Various existing recommender systems demonstrate the effectiveness of leveraging reviews in mitigating data sparsity issues (Zheng et al., 2017; Tan et al., 2016).
@@
Improving Accuracy: Improving accuracy is another challenging task for any recommender system. Recommendation accuracy largely depends on accurately learning user and item representation. Since reviews contain fine-grained opinions, utilizing them for feature learning can be beneficial. Significant research has utilized reviews to improve accuracy (Kim et al., 2016; Zhang et al., 2017; Tan et al., 2008; Seo et al., 2017). Different NLP techniques are used to extract the features from user reviews and item reviews, followed by an interaction technique to predict the ratings (Chin et al., 2018; Guan et al., 2019). For example, Chin et al. (Chin et al., 2018) utilize Word2Vec (Church, 2017) word embedding for vector representation of words for feature extraction and attention mechanism for aspect important estimation. Similarly, Guan et al. (Guan et al., 2019) use a mutual attention mechanism to capture the fine-grained interaction information by establishing correlations between user and item contextual features across multiple feature subspaces. Some studies utilize both reviews and ratings for better learning of user and item features, as well as for improving recommendation accuracy (McAuley and Leskovec, 2013).
@@
Enhancing Explainability: Explainability is a significant issue in recommender systems. A recommender system needs to be able to provide a rationale for its recommendations to a user. Traditional rating-based systems lack explainability because they solely rely on latent factor learning. Latent factors represent the hidden aspects that influence user preferences, but they usually cannot explain why a user gave a specific rating to an item. Analyzing reviews can identify the specific aspects/features that motivated users to provide certain ratings. Many models have demonstrated that utilizing reviews enhances explainability (McAuley and Leskovec, 2013; Seo et al., 2017).
@@ 4. Review-based recommender systems
The main components of recommender systems are user/item representation learning, user-item feature interaction mechanism, and overall rating or preference prediction. Traditional recommender systems take user $ID$ and item $ID$ as input and learn their features using latent factor models (Koren et al., 2009). However, these static feature-learning methods suffer from data sparsity and scalability problems. It is often beneficial to use as much information as possible, including the user review text(Chin et al., 2018), images (Yilma and Leiva, 2023), numerical rating scores (Cheng et al., 2018a), video (Mu and Wu, 2023), and audio (Bogdanov et al., 2013). Many e-commerce platforms such as Amazon, Yelp, and TripAdvisor collect user feedback as textual reviews in addition to numerical ratings. User reviews are a rich source of feedback for learning user preferences and item attributes. Because users are free to express their opinions, it is assumed that their fine-grained preferences are available in the reviews. These reviews can be utilized in various ways to solve problems in traditional rating-based recommender systems, e.g., data sparsity, and explainability problems.
Figure 2. An example of a hotel review highlighting aspects such as “service”, “location” and “cost”
Various strategies are applied to effectively utilize the reviews in recommendation systems. Based on how reviews are integrated into recommendation systems, we classify review-based recommender systems into four main categories: 1) Generic Review-based Methods, 2) Aspect-based Methods, 3) Review and Rating Fusion Methods, and 4) Ratings and Aspects Fusion Methods. Each category is discussed in details below. Additionally, relevant citations for each category are summarized in Table 1.
4.1. Generic Review-based Methods
The generic review-based method focuses on leveraging user reviews to improve recommendation accuracy. In this case, reviews are used in two different ways, a) User and Item Representation (Seo et al., 2017; Shuai et al., 2022; Almahairi et al., 2015), and b) Sentiment Analysis (Preethi et al., 2017).
a) User and Item representations: This approach primarily aims to learn comprehensive representations of both users and items from reviews, without emphasizing any particular aspect or topic. To illustrate this category, suppose a user set $U=\{u_{1},u_{2},...,u_{|U|}\}$, item set $I=\{i_{1},i_{2},...,i_{|I|}\}$, review set $E=\{e_{1},e_{2},...,e_{|E|}\}$, and an overall rating set $R=\{r_{1},r_{2},...,r_{|R|}\}$. In a traditional rating-based recommender system, user $u^{\prime}s$ preference $\hat{r}$ towards item $i$ is estimated by computing the dot product between the user vector and item vector, as shown below:
@@ 4.1. Generic Review-based Methods
where $P_{u}$ and $Q_{i}$ are the latent feature representation of user $u$ and item $i$ and $b$ is the bias. $P_{u}$ and $Q_{i}$ are learned from $U$ and $I$ respectively, e.g., $P_{u}=f(U)$ and $Q_{i}=f^{\prime}(I)$. Various algorithms can be used for the implementation of functions $f$ and $f^{\prime}$, and in some cases, they may use the same algorithm, e.g., DeepCoNN (Zheng et al., 2017)
In a generic review-based system, a user document $d_{u}$ is created by concatenating all reviews the user has written. Similarly, an item document $d_{i}$ is created by combining all reviews for that item. The features are extracted from these documents using NLP techniques (e.g., GloVe (Pennington et al., 2014), Word2Vec (Church, 2017), or BERT (Wang et al., 2024)), and latent representations of users and items are learned. For example, (Seo et al., 2017) uses attention mechanisms and convolution layers to learn user and item representations, while (Wang et al., 2018) employs a CNN for context-aware review representations. If the user and item representation learned from user reviews $d_{u}$ and item reviews $d_{i}$ are $X_{u}$ and $X_{i}$ respectively, e.g., $X_{u}=f(d_{U})$ and $X_{i}=f^{\prime}(d_{I})$, then the equation (3) for rating prediction tasks can be updated as follows:
@@
where $b$ is the bias term.
b) Sentiment Analysis: The sentiment analysis approach relies on extracting the user’s overall sentiment on the item, and the recommendations made based on sentiment score. Some studies use external sentiment analyzers, while others apply machine learning for sentiment analysis. For example, (Asani et al., 2021; Preethi et al., 2017; Chatterjee et al., 2021; Sumaia et al., 2023) use WordNet to extract user sentiment, while (Preethi et al., 2017) applies deep learning to capture sentiment directly from reviews.
4.2. Aspect-based Methods
Aspect-based methods aim to capture users’ fine-grained opinions by extracting or learning latent representations of specific aspects. Consider an aspect set $A=\{1,2,3,..|A|\}$ where $|A|$ is the number of aspects. $A_{U}$ and $A_{I}$ represent user and item aspects respectively. The aspect-based representation of user and item are $X_{u}^{A}$ and $X_{i}^{A}$ respectively. $X_{u}^{A}$ and $X_{i}^{A}$ are a function of $A_{u}$ and $A_{i}$, such that $X_{u}^{A}=f(A_{u})$ and $X_{i}^{A}=f^{\prime}(A_{i})$. For the aspect-based method, the equation (3) can be reformulated as
@@ 4.2. Aspect-based Methods
where $b$ is the biased term. The term “aspect” refers to the specific attributes of products. In Figure 2 “Location”, “service”, and “cost” represent the aspects. Naturally, a user might not be interested in all aspects of a product but one or more of them. Furthermore, the significance of an aspect to a user can differ from one item to another. Hence, it is vital to capture the aspects and their relevance to users. Aspect-aware recommender models leverage advanced NLP techniques such as topic modeling, deep neural networks, and attention mechanisms to dynamically model user preferences on various aspects.
4.3. Rating and Review Fusion Methods
In rating and review fusion methods, both rating-based and review-based features are combined to improve user and item representations. The mathematical representation for this type of method can be derived from equations (4) and (5) as follows.
(6) $\hat{r}=\zeta(P_{u},X_{u})^{\mathsf{T}}\zeta^{\prime}(Q_{i},X_{i})+b$
where $\zeta$ and $\zeta^{\prime}$ are the aggregation method integrating rating and review based features. They can be same or different for user and item. Rating and review fusion methods also extend to sentiment-focused recommendation systems. In this case, user sentiments are extracted and integrated into the user-item rating matrix to enrich the rating matrix. For example, (Osman et al., 2021) incorporates the sentiment information into the user-item rating matrix, enhancing the quality of the user-item rating matrix.
4.4. Ratings and Aspects Fusion Methods
Ratings and aspects fusion methods integrate rating-based user and item representations with aspect-based representations. The combined effect improves the accuracy of rating predictions or enhances the ranking quality of the recommendation system. Similarly, the equation for this method can be derived from (4) and (5) as follows:
(7) $\hat{r}=\zeta(P_{u},X_{u}^{A})^{\mathsf{T}}\zeta^{\prime}(Q_{i},X_{i}^{A})+b$
Different algorithms are used as aggregation functions including, machine learning, probabilistic, counterfactual, etc. For example, (Wei et al., 2024) uses contrastive learning to incorporate graph-based features from ratings to review-based features. Similarly, (Wang et al., 2023a) utilizes element-wise product as an aggregation method to integrate rating-based and review-based features.
Some other work focuses on extracting aspect-based sentiment scores and leverages them to collaborative filtering, enhancing recommendation accuracy (Hernández-Rubio et al., 2019; Al-Ghuribi et al., 2020; Osman et al., 2021; Al-Ghuribi et al., 2023a; Sumaia et al., 2023). On the other hand, (Huang et al., 2020a) computes the aspect-specific sentiment score and searches for similarity-based nearest neighbors for recommendation.
@@ 4.4. Ratings and Aspects Fusion Methods
Generic Review-based (Yang et al., 2023a; Xv et al., 2022; Choi et al., 2022; Rafailidis and Crestani, 2019; Wang et al., 2018; Seo et al., 2017; Catherine and Cohen, 2017; Zheng et al., 2017; Kim et al., 2016; Almahairi et al., 2015; Liu and Miyazaki, 2023; Wu et al., 2021; Jin et al., 2020; Osman et al., 2021; Preethi et al., 2017; Asani et al., 2021; Chatterjee et al., 2021)
Aspect-based (Li et al., 2023b; Tan et al., 2021; Xiong et al., 2021; Yang et al., 2021; Liu et al., 2021d; Ni et al., 2019; Guan et al., 2019; Chin et al., 2018; Bauman et al., 2017; He et al., 2015; Li et al., 2015; Shi et al., 2022; Liu et al., 2022; Sun et al., 2021a; Ray et al., 2020; Li and Xu, 2020; Huang et al., 2020b; Da’u et al., 2020; Da’u and Salim, 2019a)
Rating & Review Fusion (Wei et al., 2024; Chiang et al., 2023; Shuai et al., 2022; Luo et al., 2021; Zhang et al., 2021; Lyu et al., 2021; Wang et al., 2021b; Liu et al., 2021c; Gao et al., 2020b; Liu et al., 2020; Dong et al., 2020; Xia et al., 2019; Wu et al., 2019c; Liu et al., 2019a, b; Alexandridis et al., 2019; Wu et al., 2019a; Chen et al., 2019; Deng et al., 2018; Mei et al., 2018; Chen et al., 2018a; Lu et al., 2018; Chen et al., 2018b; Zhang et al., 2017; Xi et al., 2021; Wei et al., 2023; Cai et al., 2022; Du et al., 2021; Liu et al., 2021b, a; Shalom et al., 2019; Zhuang et al., 2024; Xu et al., 2018; Shoja and Tabrizi, 2019; Hung, 2020; Hu et al., 2020; Osman et al., 2021; Liao et al., 2024)
Ratings & Aspects Fusion (Wang et al., 2023b, a; Peña et al., 2020; Hyun et al., 2018; Cheng et al., 2018b, a; Tan et al., 2016; Liao et al., 2024)
As shown in Table 1, the most popular methods are Rating and Review fusion methods, followed by Aspect-based methods. The least popular method is the rating and aspect fusion method. Overall, integrating review-based features with rating-based methods is found to be used more often.
5. State-of-the-art Review-based Recommender Systems
In this section, we briefly present the state-of-the-art review-based recommender systems. Our analysis emphasizes the methodology, the rationale behind each method, features, datasets, evaluation metrics, and performance. We classify these systems into three main categories based on the methods they used: 1) Probabilistic and Topic Modeling Approaches, 2) Deep Learning-Based Approaches, and 3) Miscellaneous. This classification enables a comprehensive understanding of the diverse methods used in review-based recommender systems, ensuring that all major techniques are covered.
5.1. Probabilistic and Topic Modeling Approaches